# DRIP: DRop unImportant data Points - Enhancing Machine Learning Efficiency with Grad-CAM-Based Real-Time Data Prioritization for On-Device Training

### v1

@@ Abstract
Selecting data points for model training is critical in machine learning. Effective selection methods can reduce the labeling effort, optimize on-device training for embedded systems with limited data storage, and enhance the model performance. This paper introduces a novel algorithm that uses Grad-CAM to make online decisions about retaining or discarding data points. Optimized for embedded devices, the algorithm computes a unique DRIP Score to quantify the importance of each data point. This enables dynamic decision-making on whether a data point should be stored for potential retraining or discarded without compromising model performance. Experimental evaluations on four benchmark datasets demonstrate that our approach can match or even surpass the accuracy of models trained on the entire dataset, all while achieving storage savings of up to 39%. To our knowledge, this is the first algorithm that makes online decisions about data point retention without requiring access to the entire dataset.
@@ I Introduction
In the rapidly evolving domain of machine learning, the quantity of available data have reached unprecedented levels. While large datasets have traditionally been the bedrock of robust machine learning models, the sheer magnitude of data now available poses both opportunities and challenges. One of the primary challenges is efficient data management, especially but not only in scenarios with constrained computational and storage resources [1]. The indiscriminate data accumulation can lead to increased storage costs, longer training times, and potential overfitting due to redundant or wrong-labeled data. Moreover, in real-world applications, especially in edge computing and embedded systems [2], the availability of extensive storage and computational resources is frequently limited. Yet, new approaches do not only execute model inference on the devices but move towards so-called on-device training, which requires to store data points for model retraining in a limited storage environment. In such applications arises a need for a systematic approach to discern the utility of each collected data point and decide its retention or discard in streaming. Other applications for data selection methods target to reduce labeling effort or improving model performance [3].
@@
This raises the central research question of our study: How can selective, streaming data storage using a Grad-CAM-based scoring system improve model accuracy and reduce memory requirements for machine learning on devices in resource-constrained environments, while reducing labeling overhead without compromising model performance?
This paper delves into this challenge and presents a novel algorithm that employs Gradient-weighted Class Activation Mapping (Grad-CAM) to make informed decisions about data retention [5]. The motivation for this research stems from the increasing need for efficient data management in embedded machine learning (tinyML) [6]. Traditional data retention methods are inadequate for embedded applications, necessitating a novel approach to ensure both storage efficiency and model accuracy. Our proposed method calculates a distinct metric, termed the DRIPScore (DRIPS), which quantifies the relevance of each data point in the context of model training and performance. Using this metric, the algorithm can dynamically assess the significance of a streaming data point and make decisions about its retention.
@@ II-A Relevant Works in Data Valuation
Strumbelj and Kononenko [8] brought forth Data Shapley as a method to quantify the importance of individual data points within a dataset. By evaluating each point’s contribution to the model’s overall performance, Data Shapley offers guidance on data selection for both training and deployment. Despite its potential, Data Shapley is computationally intensive, particularly for large datasets, and operates offline, restricting streaming data importance assessments [9]. Such challenges necessitate continued research to harness its full potential in machine learning tasks.
@@ III Proposed Algorithm
Figure 1: Flowchart illustrating the seven-step process of the DRIP algorithm. The flowchart provides a visual representation of the algorithm’s sequential steps, from initial model training to the final decision on data point retention in on-device scenarios.
@@
The subsequent subsections provide a detailed breakdown of each phase, elucidating the steps involved and the rationale behind them. Fig  1 provides a graphical overview of the sequence of the DRIP algorithm
@@ III-A Training Phase
    Compute DRIPS (DRIP Score): For each image $I$ in the test dataset, compute the Grad-CAM heatmap $H(I)$. For each Grad-CAM heatmap $H(I)$, calculate the average value of the heatmap. This is done using the formula:
@@ IV-A4 Procedure
IV-B Overview of Evaluated Datasets
  1. 1.
  2. 2.
  3. 3.
  4. 4.
IV-C Neural Network Architectures Used
@@
IV-C1 Image-Based Datasets (MNIST, CIFAR-10, Plant Disease)
@@
IV-C2 Audio Dataset (Speech Commands)

### v2

# DRIP: DRop unImportant data Points - Enhancing Machine Learning Efficiency with Grad-CAM-Based Streaming Data Prioritization for On-Device Training

@@ Abstract
Selecting data points for model training is critical in machine learning. Effective selection methods can reduce the labeling effort, optimize on-device training for embedded systems with limited data storage, and enhance the model performance. This paper introduces a novel algorithm that uses Grad-CAM to make online decisions about retaining or discarding data points. Optimized for embedded devices, the algorithm computes a unique DRIP Score to quantify the importance of each data point. This enables dynamic decision-making on whether a data point should be stored for potential retraining or discarded without compromising model performance. Experimental evaluations on four benchmark datasets demonstrate that our approach can match or even surpass the accuracy of models trained on the entire dataset, while achieving storage savings of up to 39%. To our knowledge, this is the first algorithm to make online decisions about data point retention without requiring access to the entire dataset.
@@ I Introduction
In the rapidly evolving domain of machine learning, the quantity of available data have reached unprecedented levels. While large datasets have traditionally been the bedrock of robust machine learning models, the sheer magnitude of data now available poses both opportunities and challenges. One of the primary challenges is efficient data management, especially but not only in scenarios with constrained computational and storage resources [1]. The indiscriminate data accumulation can lead to increased storage costs, longer training times, and potential overfitting due to redundant or wrong-labeled data. Moreover, in many real-world applications, especially in edge computing and embedded systems [2], the storage and computational resources are limited. Yet, new approaches do not only execute model inference on the devices but move towards so-called on-device training, which requires to store data points for model retraining. In such applications arises a need for a systematic approach to discern the utility of each collected data point and decide its retention or discard in streaming is needed. Other applications for data selection methods target to reduce labeling effort or improving model performance [3].
@@
To address these challenges, we propose leveraging insights from the model’s internal decision-making process. Specifically, we use Gradient-weighted Class Activation Mapping (Grad-CAM), a technique originally developed to visualize model predictions, as the foundation for quantifying and prioritizing data points during streaming. This raises the central research question of our study: How can selective, streaming data storage using a Grad-CAM-based scoring system improve model accuracy and reduce memory requirements for machine learning on devices in resource-constrained environments, while reducing labeling overhead without compromising model performance?
This paper delves into this challenge and presents a novel algorithm that employs Grad-CAM to make informed decisions about data retention [5]. The motivation for this research stems from the increasing need for efficient data management in embedded machine learning (tinyML) [6]. Traditional data retention methods are inadequate for embedded applications, necessitating a novel approach to ensure both storage efficiency and model accuracy. Our proposed method calculates a distinct metric, termed the DRIP Score (DRIPS), which quantifies the relevance of each data point in the context of model training and performance. Using this metric, the algorithm can dynamically assess the significance of a streaming data point and make decisions about its retention.
@@ II-A Relevant Works in Data Valuation
Data valuation techniques aim to quantify the importance of individual data points for model performance, enabling more efficient training and storage. Strumbelj and Kononenko [8] brought forth Data Shapley as a method to quantify the importance of individual data points within a dataset. By evaluating each point’s contribution to the model’s overall performance, Data Shapley offers guidance on data selection for both training and deployment. Despite its potential, Data Shapley is computationally intensive, particularly for large datasets, and operates offline, restricting streaming data importance assessments [9]. Such challenges necessitate continued research to harness its full potential in machine learning tasks.
@@ III Proposed Algorithm
Figure 1: Flowchart illustrating the seven-step process of the proposed DRIP algorithm. The flowchart provides a visual representation of the algorithm’s sequential steps, from initial model training to the final decision on data point retention in on-device scenarios.
@@
The subsequent subsections provide a detailed breakdown of each phase, elucidating the steps involved and the rationale behind them. Fig. 1 provides a graphical overview of the sequence of the DRIP algorithm.
@@ III-A Training Phase
    Compute DRIP Score (DRIPS): For each image $I$ in the test dataset, compute the Grad-CAM heatmap $H(I)$. For each Grad-CAM heatmap $H(I)$, calculate the average value of the heatmap. This is done using the formula:
@@ IV-A4 Procedure
IV-B Neural Network Architectures Used
@@
IV-B1 Image-Based Datasets (MNIST, CIFAR-10, Plant Disease)
@@
IV-B2 Audio Dataset (Speech Commands)